# InComeS: Integrating Compression and Selection Mechanisms into LLMs for Efficient Model Editing

### v2

@@ Abstract
Although existing model editing methods perform well in recalling exact edit facts, they often struggle in complex scenarios that require deeper semantic understanding rather than mere knowledge regurgitation. Leveraging the strong contextual reasoning abilities of large language models (LLMs), in-context learning (ICL) becomes a promising editing method by comprehending edit information through context encoding. However, this method is constrained by the limited context window of LLMs, leading to degraded performance and efficiency as the number of edits increases. To overcome this limitation, we propose InComeS, a flexible framework that enhances LLMs’ ability to process editing contexts through explicit compression and selection mechanisms. Specifically, InComeS compresses each editing context into the key-value (KV) cache of a special gist token, enabling efficient handling of multiple edits without being restricted by the model’s context window. Furthermore, specialized cross-attention modules are added to dynamically select the most relevant information from the gist pools, enabling adaptive and effective utilization of edit information. We conduct experiments on diverse model editing benchmarks with various editing formats, and the results demonstrate the effectiveness and efficiency of our method.
@@ 1 Introduction
Model editing, also known as knowledge editing, has seen rapid progress in recent years [8, 21, 40, 46]. Its primary goal is to precisely integrate updated knowledge into a model, enabling targeted behavioral modifications while maintaining performance on unrelated tasks. Existing techniques have demonstrated strong performance in accurately recalling edited facts [43, 46, 45]. However, they often struggle in more complex editing scenarios, such as multi-hop editing composition [49, 45], natural language editing [1], and editing tasks that require reasoning and generalization [2, 46]. Moreover, recent studies [45] show that previous editing methods are prone to overfitting: they may assign excessively high probabilities to edited targets, which can distort the model’s responses to more complex or nuanced queries.
@@
To address these challenges, we introduce InComeS  [^1] https://github.com/Syon-Li/InComeS (Integrating Compression and Selection Mechanisms), a novel framework for efficient and scalable model editing. InComeS adopts context compression techniques to condense the representation of each edit into the KV cache of special gist tokens, which can be cached and reused for computational efficiency. While gisting [30] was originally developed to compress single-input prompts, we extend this approach to handle multiple edits by further introducing a specialized selection mechanism. Specifically, we augment the model with cross-attention modules that allow each input token to attend to the compressed gist representations of edits, enabling fine-grained and adaptive selection of the most relevant information. Since each edit is compressed in parallel, our framework overcomes the limitations imposed by the context window, and the specialized selection modules can be learned to enhance retrieval accuracy.
@@ 2 Preliminary
Model editing [43, 28] aims to adjust a base model $\psi$ to a post-edited model $\psi^{\prime}$ according to a set of editing information $\mathcal{T}=\{t_{1},\dots,t_{n}\}$:
$\psi^{\prime}=\text{Edit}(\psi,\{t_{1},\dots,t_{n}\})$ (1)
Here, “Edit” indicates the model editing method, while $\{t_{1},\dots,t_{n}\}$ represents the knowledge pieces to be integrated. A typical example of editing information is query-label pair $t=(x,y)$, where the goal is for the edited model to produce $y$ in response to input $x$, even if the original model does not:
$\psi(x)\neq y,\quad\psi^{\prime}(x)=y$ (2)
When the editing set contains only a single piece of information ($|\mathcal{T}|=1$), this is known as single-instance editing. In contrast, batch editing refers to the scenario where multiple pieces of knowledge are updated simultaneously ($|\mathcal{T}|>1$). Batch editing is particularly practical in real-world applications, where simultaneously updating several edits is often required. In these scenarios, it will be more efficient to integrate them into the model in a single operation.
In practice, editing information can take various forms beyond simple query-label pairs. For instance, multiple related edits can be combined to enable multihop editing, or updated knowledge may be provided as a paragraph of natural language text. In such scenarios, many traditional editing methods may struggle to produce the desired outcomes, since they are not designed to handle these diverse types of edit information. In contrast, in-context learning (ICL) approaches, where editing information is simply concatenated as contextual prefixes, offer a straightforward yet powerful solution:
$\text{Edit}_{\text{ICL}}(\psi,\{t_{1},\dots,t_{n}\})(x)=\psi(t_{1},\dots,t_{n},x)$ (3)
By leveraging the LLM’s ability to understand and reason over context, ICL can naturally accommodate a wide range of editing scenarios. Nevertheless, ICL is constrained by the context window of LLMs, and its accuracy and efficiency tend to decline when processing larger batches of edits.
@@ 3 Method
In this work, we aim to enhance the ICL-based editing approach to better understand multiple edits and accurately extract relevant information from the edit batch. Given a batch of editing information set $\mathcal{T}=\{t_{1},\dots,t_{n}\}$, an input query $x$, and the subset of its related [^2] We define related edits as those editing pieces that the model should reference when answering the query. edits $\{t_{i}|i\in\mathfrak{R}(x)\}$, we hope that our model can answer the query as effectively as a vanilla LM provided only with the relevant edits (ignoring the irrelevant editing information):
$\psi^{\prime}=\text{Edit}(\psi,\{t_{1},\dots,t_{n}\})\approx\text{Edit}(\psi,\{t_{i}|i\in\mathfrak{R}(x)\})$ (4)
@@ 3.1 Edit Compression
We adopt the concept of gisting [30], which is originally developed to compress input prompts into the representations of an extra, specially inserted token (the gist token). The condensed gist activation serves the same function as the original prompt and can be cached for later reuse, thereby improving computational and memory efficiency. While the original work primarily focuses on instruction tuning, we extend this idea to edit compression.
@@ 3.2 Edit Selection
After compressing the edit contexts, we obtain a pool of gist representations for the batch of edits. To integrate this information into the model, we introduce additional cross-attention modules that enable input tokens to attend to the edit representations. Since these representations are stored as KV caches, we leverage a similar attention mechanism to incorporate the edit information. Formally, given a token’s query state $q$, the cross-attention is computed as:
$o_{cross}=attention(q,\{gK_{0},gK_{1},gK_{2},\dots,gK_{n}\},\{gV_{0},gV_{1},gV_{2},\dots,gV_{n}\})$ (5)
We finally add the cross-attention outputs to the self-attention outputs for information aggregation.
@@
3.3 Training
Figure 2: An overview of training InComeS. The teacher model performs two forward passes: one with edit-contextualized input (1) and one with uncontextualized input (2). The cross-entropy between the outputs of (1) and (2) is used to compute a customized weight (3). The student model then compresses the edit information into KV representations using gist tokens (4). These KV caches are used to supply edit-relevant information to the query tokens (5). The final loss is computed as the sum of weighted cross entropy and KL divergence (6).
@@ 3.3 Training
$w_{x_{i}}=\max(0,\ CE(x_{i}|x_{0},\dots,x_{i-1})-CE(x_{i}|\{t_{i}|i\in\mathfrak{R}(x)\},x_{0},\dots,x_{i-1}))$ (6)
@@
In addition to token reweighting, we also adopt knowledge distillation [14] to transfer the teacher model’s knowledge about the edit information into the target model. Specifically, we apply the KL divergence to align the output distributions of the gist-contextualized student model with those of the edit-contextualized teacher model:
$KL_{x_{i}}=D_{KL}(p_{T}(x_{i}|\{t_{i}|i\in\mathfrak{R}(x)\},x_{0},\dots,x_{i-1})\ ||\ p_{S}(x_{i}|\{g_{1},\dots,g_{n}\},x_{0},\dots,x_{i-1}))$ (7)
@@
Here, $g_{1},\dots,g_{n}$ denote the cached gists for all the edits. We apply the token reweighting only to the vanilla cross-entropy term in our final loss, since we found that it would degrade effective learning of “attend-to-nothing” behavior if combined with the KL part. The explicit training details are provided in Appendix B.
@@ 4.1 Experiment setting
To verify the effectiveness of our method in complex editing scenarios, we conduct experiments on five popular datasets in model editing: the dataset for multi-hop editing MQuAKE [49], the natural language editing dataset DUNE [1], the extended version of ZsRE [43, 46], which adds a portability test set to the original ZsRE [20], and the dataset containing ripple effect samples $\text{WikiData}_{counterfact}$ [2, 46]. We report edit success rate and portability for the extend-ZsRE and $\text{WikiData}_{counterfact}$, the results for 2, 3, and 4 edits for MQuAKE [49] and new information, scientific reasoning, and debiasing for DUNE [1]. More details about the datasets and evaluation metrics can be found in the Appendix C.1 and Appendix C.2, respectively.
@@
For baselines, we select representative methods demonstrated to be powerful in relevant surveys [43, 46]. For methods that directly edit the model weights, we include ROME [24], R-ROME [10], and MEMIT [25]; for methods that adopt explicit external memory, we include SERAC [29], and IKE [48]; for methods that train additional meta-model or use implicit external memory (stores activations or nerons, etc), we adopt MEND [28], GRACE [12] and KN [3]. We also include the traditional but powerful method, like fine-tuning, LoRA [16], and ICL, which directly concatenates all the edits as the prefix context. While some similarities exist between our method and RAG, they vary considerably in problem setting and methodology. A detailed analysis is given in Appendix D.1. We choose two representative open-source models for evaluation: Llama-3.2-1B  [^4] https://huggingface.co/meta-llama/Llama-3.2-1B and Qwen2.5-7B [42]. Unless otherwise specified, we adopt an edit batch size of 100 for batch editing. More details on the baseline implementation can be found in the Appendix C.3.
@@ 4.2 Main results
We test our method on MQuAKE for the multiple-hop edit scenario, where the models are required to check multiple edits to answer each query. Because of this requirement, we mainly compare with methods designed to support batch or sequential editing. Table 1 presents our main results, which demonstrate the effectiveness of InComeS in both single-editing and batch-editing scenarios. In addition, InComeS surpasses ICL in all metrics except single 2-edits and 3-edits for Qwen2.5-7B, which shows that our method can effectively select relevant information from the editing contexts. Interestingly, single-edit specialized methods (such as ROME) collapse even in the single multi-hop query setting, revealing their incapability to handle complex editing scenarios.
Method Model Single Editing Batch Editing 2-edits 3-edits 4-edits 2-edits 3-edits 4-edits Base Llama-3.2-1B 41.79 43.51 31.58 41.79 43.51 31.58 FT-M 55.32 56.59 52.22 51.89 50.15 44.1 LoRA [16] 67.63 68.84 60.21 50.29 47.88 47.15 ROME [24] 1.97 7.77 0.55 - - - R-ROME [10] 2.72 4.29 8.53 - - - MEMIT [25] 40.21 39.54 22.34 29.71 28.87 15.7 EMMET [11] 5.4 6.61 4.74 11.49 17.16 15.89 GRACE [12] 8.45 7.13 5.24 2.03 2.68 2.06 SERAC [29] 41.84 43.51 31.58 41.84 43.45 31.74 MEND [28] 39.88 35.23 33.31 35.55 30.29 29.89 ICL 59.23 59.00 51.63 50.06 49.86 42.37 InComeS $\textbf{71.19}_{20.19\%\uparrow}$ $\textbf{72.17}_{22.32\%\uparrow}$ $\textbf{72.62}_{40.65\%\uparrow}$ $\textbf{53.93}_{7.73\%\uparrow}$ $\textbf{52.79}_{5.88\%\uparrow}$ $\textbf{52.73}_{24.45\%\uparrow}$ Base Qwen2.5-7B 44.08 44.14 30.62 44.08 44.14 30.62 FT-M 69.89 74.88 74.96 50.31 49.04 50.39 LoRA [16] 36.95 34.28 29.02 16.41 24.22 23.12 ROME [24] 9.67 7.33 8.93 - - - R-ROME [10] 10.73 6.68 3.62 - - - MEMIT [25] 44.14 46.05 32.15 43.94 46.10 30.55 EMMET [11] 26.10 38.38 26.36 40.83 45.02 33.89 GRACE [12] 15.78 13.45 13.23 5.43 7.88 4.23 SERAC [29] 55.56 59.23 53.67 42.34 40.33 39.39 MEND [28] 34.23 45.34 30.25 39.88 35.45 34.21 ICL 69.76 76.91 74.54 53.53 50.54 44.77 InComeS $66.46_{4.73\%\downarrow}$ $71.24_{7.37\%\downarrow}$ $\textbf{76.54}_{2.68\%\uparrow}$ $\textbf{55.13}_{3.00\%\uparrow}$ $\textbf{53.48}_{5.82\%\uparrow}$ $\textbf{47.91}_{7.01\%\uparrow}$
Table 1: Results on MQuAKE [49]. The difference between InComeS and ICL is marked.
@@
One of our method’s advantages is its flexibility in handling a variety of editing contexts with different formats. Unlike many traditional editing methods like ROME [24] and MEMIT [25], which require the input to follow the triplet-like fact statement format, InComeS can take edits in free-text forms without explicitly labeled subjects and objects. To verify our method’s capability for such scenarios, we adopt the DUNE dataset, which includes natural-language form edits, and the results are shown in Table 2. Following the original paper of DUNE [1], we include fine-tuning, LoRA, SERAC [29], and ICL as our baselines. The result confirms our method’s capability to handle natural language edits. Interestingly, the raw model itself is a strong baseline in the batch editing scenario, which may demonstrate the fast-evolving model capabilities over the years.
Method Model Single Editing Batch Editing New info Scientific R. Debiasing New info Scientific R. Debiasing Base Llama-3.2-1B 56.85 55.87 32.73 56.85 55.87 32.73 FT-M 57.43 53.34 35.73 57.07 53.45 33.43 LoRA 53.65 50.87 36.73 56.77 54.66 35.83 SERAC [29] 52.67 48.96 46.73 50.76 47.32 36.52 ICL 58.67 55.84 56.65 56.94 55.68 39.46 InComeS $\textbf{60.00}_{2.27\%\uparrow}$ $\textbf{58.17}_{4.17\%\uparrow}$ $54.61_{3.60\%\downarrow}$ $\textbf{57.76}_{1.44\%\uparrow}$ $\textbf{56.46}_{1.40\%\uparrow}$ $\textbf{46.14}_{16.93\%\uparrow}$ Base Qwen2.5-7B 63.44 66.03 36.95 63.44 66.03 36.95 FT-M 64.83 67.77 48.73 64.27 64.56 41.51 LoRA 63.85 63.22 39.56 62.58 65.33 43.73 SERAC [29] 64.45 63.57 33.38 56.78 58.97 31.24 ICL 65.81 65.34 35.25 66.29 66.24 40.73 InComeS $\textbf{66.83}_{1.55\%\uparrow}$ $\textbf{68.02}_{4.10\%\uparrow}$ $\textbf{62.59}_{77.56\%\uparrow}$ $65.61_{1.03\%\downarrow}$ $\textbf{67.82}_{2.39\%\uparrow}$ $\textbf{56.69}_{39.18\%\uparrow}$
Table 2: Results on DUNE [1]. The difference between InComeS and ICL is marked.
@@
We further evaluate our method on two popular editing datasets that require reasoning abilities: $Wiki_{counterfact}$ [2, 46] and the extended ZsRE [43, 46]. Table 3 shows the results. [^5] We only present methods deemed representative and powerful on these two datasets, detailed version can be found in Table 7. Our primary focus is on portability, as it serves as the most representative metric for assessing a model’s comprehensive understanding of the editing information. Overall, our method achieves performance comparable to ICL and consistently outperforms other baselines. As expected, traditional editing methods such as fine-tuning and ROME exhibit high edit success rates; however, their portability scores generally lag behind the top-performing methods, highlighting a common limitation of these approaches. In contrast, ICL-based approaches that leverage the in-context learning capabilities of LLMs demonstrate superior performance in complex editing scenarios that require reasoning, owing to the enhanced context understanding of LLMs.
Method Model $\text{WikiData}_{counterfact}$ ZsRE-extended Edit Success Portability Edit Success Portability Base Llama-3.2-1B 21.28 / 21.28 19.73 / 19.73 30.06 / 30.06 40.17 / 40.17 FT-M 97.02 / 94.58 53.43 / 47.51 99.81 / 95.94 62.80 / 54.84 LoRA [16] 98.91 / 82.61 52.87 / 43.84 99.86 / 93.18 57.43 / 44.85 ROME [24] 94.33 / 44-33 40.44 / 44-33 95.41 / 44-33 46.04 / 44-33 MEMIT [25] 44-33 / 66.94 44-33 / 23.51 44-33 / 58.79 44-33 / 25.68 MEND [28] 44-33 / 26.66 44-33 / 21.06 34-33 / 43.33 34-33 / 30.77 GRACE [12] 33.27 / 25.06 14.33 / 10.51 32.00 / 24.44 12.73 / 10.91 IKE [48] 61.70 / 34-33 45.55 / 34-33 59.15 / 43-33 57.39 / 43-33 SERAC [29] 89.56 / 78.32 60.56 / 40.45 92.69 / 89.61 66.59 / 51.61 ICL 93.31 / 82.95 65.81 / 49.75 68.86 / 60.84 62.19 / 55.58 InComeS 91.16 / 76.81 65.15 / 45.66 97.22 / 87.09 70.70 / 52.23 Base Qwen2.5-7B 22.35 / 22.35 21.46 / 21.46 36.21 / 36.21 43.86 / 43.86 FT-M 98.93 / 90.18 49.39 / 43.13 99.51 / 92.60 50.04 / 46.41 LoRA [16] 77.31 / 72.22 37.04 / 31.65 86.88 / 77.78 28.61 / 24.13 ROME [24] 92.69 / 34-33 40.25 / 34-33 97.86 / 34-33 50.43 / 34-33 MEMIT [25] 34-33 / 91.16 34-33 / 39.85 34-33 / 93.28 34-33 / 49.97 MEND [28] 34-33 / 35.13 34-33 / 15.29 34-33 / 50.91 34-33 / 38.83 GRACE [12] 31.34 / 33.77 25.60 / 18.55 33.27 / 26.79 14.35 / 11.25 IKE [48] 96.40 / 34-33 75.33 / 34-33 99.75 / 34-33 83.17 / 34-33 SERAC [29] 91.79 / 80.68 51.12 / 41.26 91.12 / 82.56 62.41 / 52.63 ICL 90.24 / 85.28 66.99 / 51.66 71.75 / 71.57 66.10 / 64.57 InComeS 90.96 / 71.44 66.69 / 47.93 97.95 / 91.29 75.63 / 61.22
Table 3: Results for $WikiData_{counterfact}$ [2, 46] and ZsRE-extended [43, 46]. The data format of each cell is in "single/batch editing results". "-" means the methods are not designed for the corresponding settings. The best two statistics are marked.
@@
We further provide a scaling-up analysis to illustrate our method’s ability to generalize to larger numbers of edits, which is the main motivation of our modification over the ICL baseline. For this analysis, we use the COUNTERFACT dataset [25], as it provides a sufficient number of editing instances. We vary the number of edits from 100 to 1000, resulting in total token counts ranging from approximately 1.2K to 12K. The results are shown in Figure 3, which shows that InComeS consistently outperforms ICL, though the base models have already been pretrained over long contexts [42]. This finding suggests that the vanilla attention mechanism alone is insufficient to effectively comprehend and precisely select the required information from the context in complex editing scenarios. In contrast, our method demonstrates greater potential for handling large-scale edits through the unified compression and selection mechanism.

### v3

@@ Abstract
Although existing model editing methods perform well in recalling exact edit facts, they often struggle in complex scenarios that require deeper semantic understanding rather than mere knowledge regurgitation. Leveraging the strong contextual reasoning abilities of large language models (LLMs), in-context learning (ICL) becomes a promising editing method by comprehending edit information through context encoding. However, this method is constrained by the limited context window of LLMs, leading to degraded performance and efficiency as the number of edits increases. To overcome this limitation, we propose InComeS, a flexible framework that enhances LLMs’ ability to process editing contexts through explicit compression and selection mechanisms. Specifically, InComeS compresses each editing context into the key-value (KV) cache of a special token, enabling efficient handling of multiple edits without being restricted by the model’s context window. Furthermore, specialized cross-attention modules are added to dynamically select the most relevant information from the pool of special tokens, enabling adaptive and effective utilization of edit information. We conduct experiments on diverse model editing benchmarks with various editing formats, and the results demonstrate the effectiveness and efficiency of our method. Corresponding author.
InComeS: Integrating Compression and Selection Mechanisms into LLMs for Efficient Model Editing
Shuaiyi Li1thanks: Work done during internship at Tencent AI Lab, Zhisong Zhang2,†, Yang Deng4, Chenlong Deng3, Tianqing Fang3, Hongming Zhang3, Haitao Mi3, Dong Yu3, Wai Lam1,† 1The Chinese University of Hong Kong, 2City University of Hong Kong, 3Tencent AI Lab, 4Singapore Management University {sli, wlam}@se.cuhk.edu.hk, zhisong.zhang@cityu.edu.hk
@@ 1 Introduction
Model editing, also known as knowledge editing, has seen rapid progress in recent years Fang et al. (2024); Li et al. (2024); Wang et al. (2024); Zhang et al. (2024a). Its primary goal is to precisely integrate updated knowledge into a model, enabling targeted behavioral modifications while maintaining performance on unrelated tasks. Existing techniques have demonstrated strong performance in accurately recalling edited facts Yao et al. (2023); Zhang et al. (2024a, 2025a). However, they often struggle in more complex editing scenarios, such as multi-hop editing composition Zhong et al. (2023b); Zhang et al. (2025a), natural language editing Akyürek et al. (2023), and editing tasks that require reasoning and generalization Cohen et al. (2024); Zhang et al. (2024a). Moreover, recent studies Zhang et al. (2025a) show that previous editing methods are prone to overfitting: they may assign excessively high probabilities to edited targets, which can distort the model’s responses to more complex or nuanced queries.
@@
To address these challenges, we introduce InComeS  [^1] https://github.com/Syon-Li/InComeS (Integrating Compression and Selection Mechanisms), a novel framework for efficient and scalable model editing. InComeS adopts context compression techniques to condense the representation of each edit into the KV cache of special gist tokens, which can be cached and reused for computational efficiency. While gisting Mu et al. (2023) was originally developed to compress single-input prompts, we extend this approach to handle multiple edits by further introducing a specialized selection mechanism. We further augment the model with cross-attention modules that allow each input token to attend to the compressed gist representations of edits, enabling fine-grained and adaptive selection of the most relevant information. Since each edit is compressed in parallel, our framework overcomes the limitations imposed by the context window, and the specialized selection modules can be learned to enhance retrieval accuracy.
@@ 2 Preliminary
Model editing Yao et al. (2023); Mitchell et al. (2022a) aims to adjust a base model $\psi$ to a post-edited model $\psi^{\prime}$ according to a set of editing information $\mathcal{T}=\{t_{1},\dots,t_{n}\}$: $\psi^{\prime}=\text{Edit}(\psi,\{t_{1},\dots,t_{n}\})$. Here, “Edit” indicates the model editing method, while $\{t_{1},\dots,t_{n}\}$ represents the knowledge pieces to be integrated. A typical example of editing information is query-label pair $t=(x,y)$, where the goal is for the edited model to produce $y$ in response to input $x$, even if the original model does not: $\psi(x)\neq y,\psi^{\prime}(x)=y$. When the editing set contains only a single piece of information ($|\mathcal{T}|=1$), this is known as single-instance editing. In contrast, batch editing refers to the scenario where multiple pieces of knowledge are updated simultaneously ($|\mathcal{T}|>1$). Batch editing is particularly practical in real-world applications, where simultaneously updating several edits is often required. In these scenarios, it will be more efficient to integrate them into the model in a single operation.
In practice, editing information can take various forms beyond simple query-label pairs. For instance, multiple related edits can be combined to enable multihop editing, or updated knowledge may be provided as a paragraph of natural language text. In such scenarios, many traditional editing methods may struggle to produce the desired outcomes, since they are not designed to handle these diverse types of editing information. In contrast, in-context learning (ICL) approaches, where editing information is simply concatenated as contextual prefixes, offer a straightforward yet powerful solution: $\text{Edit}_{\text{ICL}}(\psi,\{t_{1},\dots,t_{n}\})(x)=\psi(t_{1},\dots,t_{n},x)$. By leveraging the LLM’s ability to understand and reason over context, ICL can naturally accommodate a wide range of editing scenarios. Nevertheless, ICL is constrained by the context window of LLMs, and its accuracy and efficiency tend to decline when processing larger batches of edits.
@@ 3 Method
In this work, we aim to enhance the ICL-based editing approach to better understand multiple edits and accurately extract relevant information from the edit batch. Given a batch of editing information set $\mathcal{T}=\{t_{1},\dots,t_{n}\}$, an input query $x$, and the subset of its related [^2] We define related edits as those editing pieces that the model should reference when answering the query. edits $\{t_{i}|i\in\mathfrak{R}(x)\}$, we hope that our model can answer the query as effectively as a vanilla LM provided only with the relevant edits (ignoring the irrelevant editing information): $\psi^{\prime}=\text{Edit}(\psi,\{t_{1},\dots,t_{n}\})\approx\text{Edit}(\psi,\{t_{i}|i\in\mathfrak{R}(x)\})$.
@@ 3.1 Edit Compression
We adopt the concept of gisting Mu et al. (2023), which is originally developed to compress input prompts into the representations of an extra, specially inserted token (the gist token). The condensed gist activation serves the same function as the original prompt and can be cached for later reuse, thereby improving computational and memory efficiency. While the original work primarily focuses on instruction tuning, we extend this idea to edit compression.
@@ 3.2 Edit Selection
After compressing the edit contexts, we obtain a pool of gist representations for the batch of edits. To integrate this information into the model, we introduce additional cross-attention modules that enable input tokens to attend to the edit representations. Since these representations are stored as KV caches, we leverage a similar attention mechanism to incorporate the edit information. Formally, given a token’s query state $q$, the cross-attention is computed as: $o_{cross}=attention(q,\{gK_{0},gK_{1},gK_{2},\dots,gK_{n}\},\allowbreak\{gV_{0},gV_{1},gV_{2},\dots,gV_{n}\})$. We finally add the cross-attention outputs to the self-attention outputs for information aggregation.
@@
3.3 Meta Training
Figure 2: An overview of the one-time meta training of InComeS. The teacher model performs two forward passes: one with edit-contextualized input (1) and one with uncontextualized input (2). The cross-entropy between the outputs of (1) and (2) is used to compute a customized weight (3). The student model then compresses the edit information into KV representations using gist tokens (4). These KV caches are used to supply edit-relevant information to the query tokens (5). The final loss is computed as the sum of weighted cross entropy and KL divergence (6).
@@ 3.3 Meta Training
$\begin{split}w_{x_{i}}=&\max(0,\ CE(x_{i}|x_{0},\dots,x_{i-1})-\\ &CE(x_{i}|\{t_{i}|i\in\mathfrak{R}(x)\},x_{0},\dots,x_{i-1}))\end{split}$ (1)
@@
In addition to token reweighting, we also adopt knowledge distillation Hinton et al. (2015) to transfer the teacher model’s knowledge about the edit information into the target model. Specifically, we apply the KL divergence to align the output distributions of the gist-contextualized student model with those of the edit-contextualized teacher model:
$\begin{split}KL_{x_{i}}=&D_{KL}(p_{T}(x_{i}|\{t_{i}|i\in\mathfrak{R}(x)\},x_{0},\dots,x_{i-1})\ ||\ \\ &p_{S}(x_{i}|\{g_{1},\dots,g_{n}\},x_{0},\dots,x_{i-1}))\end{split}$ (2)
@@
Here, $g_{1},\dots,g_{n}$ denote the cached gists for all the edits. We apply the token reweighting only to the vanilla cross-entropy term in our final loss, since we found that it would degrade effective learning of "attend-to-nothing" behavior if combined with the KL part. The explicit training details are provided in Appendix A.
Method Single Editing Batch Editing Llama-3.2-1B Base 38.96 38.96 FT-M 54.71 48.71 LoRA 65.56 48.44 ROME 3.43 - R-ROME 5.18 - MEMIT 34.03 24.76 EMMET 5.58 14.85 GRACE 6.94 2.26 SERAC 38.97 39.01 MEND 36.41 31.91 RECIPE 55.49 45.81 DR-IKE 49.44 38.67 ICL 56.62 47.43 InComeS 71.99 53.15 Qwen2.5-7B Base 39.61 39.61 FT-M 73.24 49.91 LoRA 33.42 21.25 ROME 8.64 - R-ROME 7.01 - MEMIT 40.78 40.20 EMMET 30.28 39.91 GRACE 14.15 5.85 SERAC 56.15 40.69 MEND 36.61 36.51 RECIPE 58.55 46.58 DR-IKE 55.61 41.37 ICL 73.74 49.61 InComeS 71.41 52.17
Table 1: Results on MQuAKE Zhong et al. (2023b). The full version can be found in Table 9. More analysis about the performance gain of different model scale can be found in Appendix C.9.
Method Single Editing Batch Editing Llama-3.2-1B Base 48.48 48.48 FT-M 48.83 47.89 LoRA 47.08 49.08 SERAC 49.45 44.87 ICL 57.05 50.69 InComeS 57.59 53.45 Qwen2.5-7B Base 55.47 55.47 FT-M 60.44 56.78 LoRA 55.54 57.21 SERAC 53.80 48.99 ICL 55.47 57.75 InComeS 65.81 63.37
Table 2: Results on DUNE Akyürek et al. (2023). Detailed version can be found in Table 10.
Method WikiDatacounterfact ZsRE-extended Single Batch Single Batch Llama-3.2-1B Base 19.73 19.73 40.17 40.17 FT-M 53.43 47.51 62.80 54.84 LoRA 52.87 43.84 57.43 44.85 ROME 40.44 - 46.04 - MEMIT 46.34 23.51 43.26 25.68 MEND 24.98 21.06 37.53 30.77 GRACE 14.33 10.51 12.73 10.91 IKE 45.55 - 57.39 - SERAC 60.56 40.45 66.59 51.61 ICL 65.81 44.75 62.19 51.58 InComeS 65.15 45.66 70.70 52.23 Qwen2.5-7B Base 21.46 21.46 43.86 43.86 FT-M 49.39 43.13 50.04 46.41 LoRA 37.04 31.65 28.61 24.13 ROME 40.25 - 50.43 - MEMIT 43.58 39.85 53.23 49.97 MEND 20.45 15.29 43.26 38.83 GRACE 25.60 18.55 14.35 11.25 IKE 65.33 - 70.17 - SERAC 51.12 41.26 62.41 52.63 ICL 66.99 51.66 66.10 60.57 InComeS 66.69 47.93 75.63 61.22
Table 3: Portability results on WikiDatacounterfact Cohen et al. (2024); Zhang et al. (2024a) and ZsRE-extended Zhang et al. (2024a); Yao et al. (2023). Full results can be found at Table 7.
@@ 4.1 Experiment setting
To verify the effectiveness of our method in complex editing scenarios, we conduct experiments on five popular datasets in model editing: the dataset for multi-hop editing MQuAKE Zhong et al. (2023b), the natural language editing dataset DUNE Akyürek et al. (2023), the extended version of ZsRE Yao et al. (2023); Zhang et al. (2024a), which adds a portability test set to the original ZsRE Levy et al. (2017), and the dataset containing ripple effect samples $\text{WikiData}_{counterfact}$ Cohen et al. (2024); Zhang et al. (2024a). We report edit success rate and portability for the extend-ZsRE and $\text{WikiData}_{counterfact}$, the results for 2, 3, and 4 edits for MQuAKE Zhong et al. (2023b) and new information, scientific reasoning, and debiasing for DUNE Akyürek et al. (2023). More details about the datasets and evaluation metrics can be found in the Appendix B.1 and Appendix B.2, respectively.
@@
For baselines, we select representative methods demonstrated to be powerful in relevant surveys Yao et al. (2023); Zhang et al. (2024a). For methods that directly edit the model weights, we include ROME Meng et al. (2022), R-ROME Gupta et al. (2024a), and MEMIT Meng et al. (2023); for methods that adopt explicit external memory, we include SERAC Mitchell et al. (2022b), IKE Zheng et al. (2023), and DR-IKE Nafee et al. (2025); for methods that train additional meta-model or use implicit external memory (stores activations or nerons, etc), we adopt MEND Mitchell et al. (2022a), GRACE Hartvigsen et al. (2022), KN Dai et al. (2022), and RECIPE Chen et al. (2024). We also include the traditional but powerful method, like fine-tuning, LoRA Hu et al. (2022), and ICL, which directly concatenates all the edits as the prefix context. While some similarities exist between our method and RAG, they vary considerably in problem setting and methodology. A detailed analysis is given in Appendix C.1. We choose two representative open-source models for evaluation: Llama-3.2-1B  [^4] https://huggingface.co/meta-llama/Llama-3.2-1B and Qwen2.5-7B Yang et al. (2024) (More results about Qwen3-8B-base is in Appendix C.10). Unless otherwise specified, we adopt an edit batch size of 100 for batch editing. More details on the baseline implementation can be found in the Appendix B.3.
@@ 4.2 Main results
We test our method on MQuAKE for the multiple-hop edit scenario, where the models are required to check multiple edits to answer each query. Because of this requirement, we mainly compare with methods designed to support batch or sequential editing. Table 1 presents our main results, which demonstrate the effectiveness of InComeS in both single-editing and batch-editing scenarios. In addition, InComeS surpasses ICL in all metrics except the single editing setting for Qwen2.5-7B, which shows that our method can effectively select relevant information from the editing contexts. Interestingly, single-edit specialized methods (such as ROME) collapse even in the single multi-hop query setting, revealing their incapability to handle complex editing scenarios. To verify the effectiveness of our method on different model scales, we conduct further analysis on the performance gain of different model scales in Appendix C.9.
@@
One of our method’s advantages is its flexibility in handling a variety of editing contexts with different formats. Unlike many traditional editing methods like ROME Meng et al. (2022) and MEMIT Meng et al. (2023), which require the input to follow the triplet-like fact statement format, InComeS can take edits in free-text forms without explicitly labeled subjects and objects. To verify our method’s capability for such scenarios, we adopt the DUNE dataset, which includes natural-language form edits, and the results are shown in Table 2. Following the original paper of DUNE Akyürek et al. (2023), we include fine-tuning, LoRA, SERAC Mitchell et al. (2022b), and ICL as our baselines. The result confirms our method’s capability to handle natural language edits. Interestingly, the raw model itself is a strong baseline in the batch editing scenario, which may demonstrate the fast-evolving model capabilities over the years.
@@
We further evaluate our method on two popular editing datasets that require reasoning abilities: $Wiki_{counterfact}$ Cohen et al. (2024); Zhang et al. (2024a) and the extended ZsRE Yao et al. (2023); Zhang et al. (2024a). Table 3 shows the results. Our primary focus is on portability, as it serves as the most representative metric for assessing a model’s comprehensive understanding of the editing information. Overall, our method achieves performance comparable to ICL and consistently outperforms other baselines. More analysis about the detailed results is given in Appendix C.8.
@@
We further provide a scaling-up analysis to illustrate our method’s ability to generalize to larger numbers of edits, which is the main motivation of our modification over the ICL baseline. For this analysis, we use the COUNTERFACT dataset Meng et al. (2023), as it provides a sufficient number of editing instances. We vary the number of edits from 100 to 1000, resulting in total token counts ranging from approximately $1.2K$ to $12K$. The results are shown in Figure 3, which shows that InComeS consistently outperforms ICL, though the base models have already been pretrained over long contexts Yang et al. (2024). This finding suggests that the vanilla attention mechanism alone is insufficient to effectively comprehend and precisely select the required information from the context in complex editing scenarios. In contrast, our method demonstrates greater potential for handling large-scale edits through the unified compression and selection mechanism.